OpenStudy (iwanttogotostanford):

What is the simplified form of the quantity of x plus 6, all over the quantity of x plus 4 + the quantity of x minus 3, all over 3? the quantity of x squared plus 3x minus 18, all over 3 times the quantity of x plus 4 the quantity of x squared plus 4x plus 6, all over the quantity of x plus 7 the quantity of x squared plus 3x minus 18, all over the quantity x plus 7 the quantity of x squared plus 4x plus 6, all over 3 times the quantity of x plus 4

OpenStudy (iwanttogotostanford):

@Cardinal_Carlo I will use the equation maker this time!

OpenStudy (iwanttogotostanford):

\[x ^{2}+4x+6 \over 3(x+4)\]

OpenStudy (iwanttogotostanford):

@Cardinal_Carlo I did it haha

OpenStudy (anonymous):

So you have to solve for x right?

OpenStudy (iwanttogotostanford):

that is my answer

OpenStudy (iwanttogotostanford):

yes

OpenStudy (anonymous):

good job @iwanttogotostanford

OpenStudy (iwanttogotostanford):

thank you! I finally got it right

OpenStudy (iwanttogotostanford):

i thought i didn't but i did

OpenStudy (anonymous):

So first open the brackets. \[ \frac{ x^2+4x+6 }{ 3x+12 }\]

OpenStudy (anonymous):

Just double-checking our work: \[\frac{ x + 6 }{ x + 4 } + \frac{ x - 3 }{ 3 }\] \[\frac{3 \left( x + 6 \right) + \left( x +4 \right) \left( x - 3 \right) }{ 3\left( x + 4 \right) }\] \[\frac{ x ^{2} + 4x + 6 }{ 3x + 12 }\] Yay!

OpenStudy (iwanttogotostanford):

ok, thanks!

OpenStudy (iwanttogotostanford):

i need help studying for my math test and have a couple of questions, can i still have help?

OpenStudy (anonymous):

You got this concept on lock @iwanttogotostanford. Great job!

OpenStudy (anonymous):

Sure just post them up, and close this one.